OpenStudy (anonymous):

find the missing number in the proportion. 5/6 = x/12

OpenStudy (anonymous):

cross-multiply, you get 6x = 60 Divide both sides by 6, you get x = 10

OpenStudy (anonymous):

set up your proportion 5/6 =x/12 cross multiply to get 60=6x then divide both sides by 6 then solve for x to get 10